An aqueous solution containing 0.01 g KIO3 (formula weight = 214.0) was treated with an excess of KI solution. The solution was acidified with HCI. The liberated I2 consumed 45.0 mL of thiosulphate solution decolourise the blue starch-iodine complex. Calculate the molarity of the sodium thiosulphate solution.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. The redox reaction involved are : 〖IO〗_3^- + 5I- + 6H+ → 3I2 + 3H2O I2 + 2S2O_3^(2-) → 2I- + S4O_6^(2-) Millimol of KIO3 used = 0.1/214 × 1000 = 0.467 ⇒ millimol of I2 formed = 3 × 0.467 = 1.4 ⇒ millimol of Na2S2O3 consumed = 2 × 1.4 = 2.8 ⇒ Molarity of Na2S2O3 = 2.8/45 = 0.062 M